OpenStudy (anonymous):

What is the product of 12t/t-3 * 5t - 5/16t^3

OpenStudy (compassionate):

Can you write it using the equation editor?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it \[\huge \frac{ 60t^2 }{ t-3}-\frac{ 5 }{ 4096t^3 }\]

OpenStudy (anonymous):

\[\frac{ 12t }{ t-3 } \times \frac{ 5t - 15 }{ 16t^3 }\]

OpenStudy (anonymous):

this is the equestion @vetdoc ??

OpenStudy (anonymous):

\[\huge \frac{ 3(5t-15) }{ 4(t-3)t^2 }\]

OpenStudy (anonymous):

The question is what I typed.

OpenStudy (anonymous):

no iam asking is this the answer!!

OpenStudy (anonymous):

That's not one of the choices.

OpenStudy (anonymous):

expand my answer then!

OpenStudy (anonymous):

\[\frac{15}{4 \times t ^{2}}\]

OpenStudy (anonymous):

it's this

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

\[\huge \frac{ 15 }{ 4t^2 }\]